OpenStudy (anonymous):

(3a-4b+2c)-(4a-5c+2b) please simplify this polynomial @midhun.madhu1987 can u help w/ this?

OpenStudy (anonymous):

or anyone..

OpenStudy (dangerousjesse):

Simplify the following: \[3 a-4 b+2 c-(4 a-5 c+2 b) \]Distribute -1 over \( 4 a+2 b-5 c\). \[-(4 a+2 b-5 c) = -4 a-2 b+5 c: \]\[3 a-4 b+2 c+-4 a-2 b+5 c \]Group like terms in \(3 a-4 b+2 c-4 a-2 b+5 c.\) Grouping like terms, \(3 a-4 b+2 c-4 a-2 b+5 c = (2 c+5 c)+(-4 b-2 b)+(3 a-4 a):\) \[(2 c+5 c)+(-4 b-2 b)+(3 a-4 a) \]Add like terms in \(2 c+5 c.\) \[2 c+5 c = 7 c: \]\[7 c+(-4 b-2 b)+(3 a-4 a) \]Combine like terms in \(-4 b-2 b.\) \[-4 b-2 b = -6 b: \]\[7 c+-6 b+(3 a-4 a) \]Combine like terms in \(3 a-4 a.\) \[3 a-4 a = -a: \]

OpenStudy (dangerousjesse):

Which leaves you with...?

OpenStudy (anonymous):

thank you, i get it now, can you help me with another? @DangerousJesse

OpenStudy (dangerousjesse):

Sure

OpenStudy (anonymous):

1)Factor each expression completely. x^2-5x-14

OpenStudy (anonymous):

2)Factor each expression completely. m^2-100

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

Factor the following: \[x^2-5 x-14 \]Factor \(x^2-5 x-14\) by finding two numbers whose product is \(-14\) and whose sum is \(-5\). The factors of \(-14\) that sum to \(-5\) are \(2\) and \(-7\). So, \(x^2-5 x-14 = (x+2) (x-7):\)

OpenStudy (dangerousjesse):

So, your answer is?

OpenStudy (anonymous):

(-2 + -1x)(7 + -1x)

OpenStudy (dangerousjesse):

\[(x+2)(x-7)\]

OpenStudy (anonymous):

thanks

OpenStudy (dangerousjesse):

No problem :)